# Optimal Selection Using Algorithmic Rankings with Side Information

### v2

@@ Abstract
Motivated by online platforms such as job markets, we study an agent choosing from a list of candidates, each with a hidden quality that determines match value. The agent observes only a noisy ranking of the candidates plus a binary signal that indicates whether each candidate is ”free” or ”busy.” Being busy is positively correlated with higher quality, but can also reduce value due to decreased availability. We study the agent’s optimal selection problem in the presence of ranking noise and free-busy signals and ask how the accuracy of the ranking tool impacts outcomes. In a setting with one high-valued candidate and an arbitrary number of low-valued candidates, we show that increased accuracy of the ranking tool can result in reduced social welfare. This can occur for two reasons: agents may be more likely to make offers to busy candidates, and (paradoxically) may be more likely to select lower-ranked candidates when rankings are more indicative of quality. We further discuss conditions under which these results extend to more general settings.
@@ 1 Introduction
Consider the following setting: you are a hiring manager, trying to find a candidate to interview for a hiring slot. You know that some candidates are better suited for your role than others - ideally, you would like to interview the best candidate available. To assist you with this goal, you might obtain a (noisy) ranking of the candidates from a commercial employment site, such as LinkedIn, Indeed, or other ranking tools. If the ranking is “good” (if the expected utility of candidates decreases as you navigate further down the list), the best action for the employer is clear: always pick the top-ranked candidate. Also, as the ranking tool becomes more accurate, the employer is only more happy with their choice, as the expected value of the top-ranked candidate only increases.
However, in many cases, employers have access to side information, besides the ranking. For example, a hiring manager may know that some candidates are less likely to accept any eventual offer because they is already employed (see filippas2023advertising for an empirical example of such a setting). We call such a candidate “busy”, as compared to a “free” candidate without these competing sources of employment. If a busy candidate would never accept an offer, then the choice is clear: the best option is to select the highest-ranked free candidate. However, if a busy candidate still has some chance of accepting an offer, the possible strategy space changes. In many markets there is an opposing force to the penalty for being busy: it may be the case that high-value candidates are more likely to be busy, and thus being busy is a signal of candidate quality. For example, this could be because high value candidates are more likely to have already been identified and hired by competing firms [^1] Throughout this paper, we will make the assumption that some candidates are better fits for the job than others (e.g. higher or lower value). We recognize that this assumption drastically simplifies the multi-faced strengths and weaknesses each candidate may bring, and discuss this (and other) assumptions in more depth in Section 2.. In many settings it may not be possible to resolve the uncertainty inherent in a busy signal: for example, the candidate themselves may not know whether or not they would be willing to move without going through the interview process, at which point the firm has already used up its interview slot.
@@
Moreover, a firm’s hiring strategy has impacts on the rest of society. For example, consider that a firm’s choice to interview a busy (already employed) candidate over a free (unemployed) candidate may be more or less valuable to the candidate. An unemployed candidate may have higher utility for an interview than an unemployed candidate. Alternatively, an employed candidate may have a stronger ability to negotiate a better offer, which could increase their welfare more. While we do not directly model such externalities, we will use “probability of picking a busy candidate” and other such quantities as benchmarks to understand the impact of a ranking tool on broader society. Specifically, we will be interested in how these quantities change as the accuracy of the ranking tool increases. Do these measures uniformly improve or degrade as the firm uses a more accurate ranking tool?
@@
Motivation: More broadly, in this paper our goal is to study a stylized model of an agent using an algorithmic tool to make a decision, and combining the information encoded by the tool with her own source of information. Given the importance attached to continually increasing the accuracy of algorithmic tools, we wish to show that when these tools are used as part of a broader system involving human discretion and alternative sources of information, the returns to increased algorithmic accuracy are less clear: in particular, relevant metrics of social welfare could deteriorate under increased accuracy. We hope that our results will lead to further research in other scenarios where algorithmic tools are used alongside other sources of information, and in particular into what properties of such algorithmic tools are helpful for societal welfare. While we will use the employer hiring job candidates as our motivating example, we note that this setting is quite broad and could model a range of other scenarios where humans make decisions. For example, consider a hungry customer navigating a ranking of restaurants, each “crowded” or “not crowded”: a crowded restaurant is more likely to be of good, but also has a longer wait.
@@
Summary: The rest of this paper is as follows: in Section 2, we will present our model and assumptions, as well as a permutation model that we believe may be of broader interest. In Section 3, we discuss the connection of our paper to related work, including the extensive literature on herding and the Pandora’s Box problem. Next, Section 4 analyzes the optimal strategy that a firm navigating a ranking might take: should she pick the first busy candidate, the first free candidate, or some other candidate entirely? In Section 5 we study notions of social welfare, like how frequently firms preferentially select job candidates who are already employed (potentially modeling unemployment discrimination) or avoid doing so (potentially modeling collusion in job hiring). We also analyze how often firms select the top ranked candidate, a measure of how incentive compatible the ranking is with potential implications for the welfare of the company creating the ranking tool. Throughout, we consider how increasing the accuracy of the ranking tool can affect the strategy, welfare, and fairness results in previous sections. Surprisingly, we show that increased accuracy can reduce measures of social welfare. In Section 6 we discuss extensions of our model and in Section 7 we conclude by discussing implications of our results. All proofs are deferred to the appendix.
@@ 2 Model, notation, and assumptions
Model and notation: First, we present our model. There are $N$ candidates, of which the agent (alternatively, firm) wishes to select exactly one: for example, to interview a job candidate. The candidates have different true values to the firm, where $v_{i}$ denotes the value of candidate $i$, though the firm cannot know the true value of candidates at each ranking (this only holds in the limit if a perfectly noiseless algorithmic ranking tool). The goal of the firm is to maximize their expected utility. In order to help with this decision, the firm has access to a ranking tool which produces noisy permutations over the candidates: $\sigma\sim\mathcal{P}$, where $\sigma_{i}$ denotes the index of the candidate ranked $i$th in permutation $\sigma$ with value $v_{\sigma_{i}}$. We will require that the expected value of candidates decreases in the ranking ($\mathbb{E}_{\sigma\sim\mathcal{P}}[v_{\sigma_{i}}]$ is decreasing in $i$).
@@
Permutation distribution: When comparing two distributions $\mathcal{P}$ and $\mathcal{P}^{\prime}$, we will use notation $P[\sigma]$ and $P[\sigma]^{\prime}$ to denote the probability $\mathcal{P}$ and $\mathcal{P}^{\prime}$ place on $\sigma$, respectively. In general, we will assume that the probability of observing each permutation is drawn from a Plackett-Luce distribution plackett1975analysis; luce1959individual). This commonly-used model is where i.i.d. noise is added to the true value of each item to produce $\hat{}v_{i}=v_{i}+\epsilon$, for $\epsilon$ drawn from a Gumbel distribution: then, the items are sorted in order of the noised values $\{\hat{}v_{i}\}$. Throughout, when we say “increased accuracy”, we will mean a smaller degree of noise (smaller standard deviation) in the Gumbel distribution. However, many of our results rely only weakly on the Plackett-Luce model, and in the relevant appendices we state the sufficient properties on the permutation distribution in order for our results to apply.
@@
Firm strategy: In addition to the ranking, the firm has access to a single bit of side information about each candidate: specifically, a status vector $s$, where $s_{i}=1$ if the candidate ranked in position $i$ is free, and $s_{i}=0$ if the candidate is busy. We assume that higher-valued candidates are less likely to be available, in particular, that $v_{i}\geq v_{j}$ implies $p_{i}\leq p_{j}$, where $p_{i}$ gives the probability that the $i$th candidate is free. We also assume that each candidate’s probability of being free or busy is independent of every other candidate: specifically, each free/busy realization is drawn from a Bernoulli distribution $s_{i}\sim Bern(p_{i})$ before the firm observes the ranking. Throughout this paper we will take these probabilities as fixed and assume each candidate reports her free-busy status honestly: in the conclusion we will discuss extensions of this model.
@@ 3 Related work
Our model has connections to celebrated models of agents strategically picking items, such as the hiring problem and Pandora’s box problem krengel1977semiamarts; weitzman1978optimal. However, our setting has key differences from these. Foremost among them is the signaling mechanism: rather than each item having an independent, known distribution of value, we model the scenario where the free/busy signal is generated by items with different probabilities, given different values. This structure, in conjunction with the ranking, means that the free/busy status of item $i$ can affect the expected value of item $j$ – generally not captured in existing models. Furthermore, in our setting the agent must pick a single item and commit to it without being able to observe its true value. This captures the setting where it is difficult or impossible to evaluate the item without committing to it first – for example, in hiring, when a candidate’s true “value” may not become apparent until after weeks or months of work.
@@
Additionally, our work has connections to long literature on herding (also known as information cascades) banerjee1992simple; bikhchandani1992theory; welch1992sequential, a pattern where it may be optimal for a decision-maker to follow actions taken by previous people, even ignoring their own information. Such patterns can lead to “cascades” where multiple sequential decision-makers each follow actions taken by previous decision-makers, rather than following their own information. In our setting, when an agent chooses to pick a lower ranked busy candidate over a higher ranked free, this could be seen as herding, since the free/busy status of candidates could be seen as created by actions taken by previous decision-makers, whereas the ranking is assumed to be private (personal) information. To our knowledge, herding has not been extensively studied in ranking settings: our model interestingly allows for herding to be moderated by the strength of private information through the quality of the ranking and the index of the first busy candidate.
@@
Our paper also has connections to the literature on algorithmic monoculture kleinberg2021algorithmic, which studies when utility may be decreased when two agents rely on the same algorithmic ranking, rather than their own (uncorrelated) rankings. Variants of this setting have been explored, such as how algorithmic monoculture can lead to individuals experiencing correlated outcomes (“outcome homogenization”) NEURIPS2022_17a234c9, or the implications of monoculture and of noisy matching in two-sided matching markets. For example, peng2023monoculture; peng2024wisdom; baek2025hiring all model settings where firms rely on common algorithmic tools to make hiring decisions, and the potential pitfalls of such an approach.
@@
Common features include the model of multiple agents taking strategic actions in relation to imperfect rankings, as well as the notion of a penalty for picking an item that another agent has picked. Key differences in our model include the fact that this penalty is softened – in particular, an agent can derive non-zero utility for picking an item that another has already chosen. This dramatically complicates the strategy space. Multiple works have considered the question of fairness in ranking - zehlike2021fairness offers a helpful survey. In particular, NEURIPS2021_63c3ddcc describes a notion of fairness relating to how often items are presented in the top $k$, given that they have some probability of truly being the best $k$ items (relating to our notion of an item being “picked”). Issues of fairness and strategy have also been studied in the related area of algorithmic monoculture blackraghavanbarocas22; jain2023algorithmic; cooper2023my; NEURIPS2022_ba4caa85, while peng2023reconciling studies the tradeoff between diversity and accuracy, showing that consumption constraints (e.g. only top item can be consumed) explains away a tension between these two goals.
@@
Finally, there are multiple papers (empirical and theoretical) exploring models of firms using prediction tools to compete with each other. For example, filippas2023advertising explicitly studies an empirical version of this where job candidates could pay to send a signal of availability to potential employers. Intriguingly, in this setting it seemed like such a signal did not lead employers to preferentially pick busy candidates, with most employers choosing a “first-free” strategy, perhaps indicating a parameter regime where this strategy was typically dominant. Michellespaper explores this setting in a dynamical pricing model, giving conditions where pricing levels for such a signal induce a truthful response. Additionally, jagadeesan2023improved explores a setting where improved data representation (as modeled by reduced Bayes risk) can paradoxically increase the error that users experience when they choose between two competing firms. 10.1145/3490486.3538364 studies statistical discrimination in stable matchings where candidates have preferences over firms, who only have noisy access to signals about candidate quality, and castera2024correlation studies a setting where different candidates have differing degrees of correlation in their predictions across models, which leads to tensions between societal welfare and individual candidate welfare.
@@ 4 Superstar: Optimal Selection Strategies
1Define $R=\frac{v_{1}/\gamma_{1}-v_{2}}{v_{1}-v_{2}/\gamma_{2}}\cdot r$.
2 if $R\leq 1$ then
3  Calculate $j^{*}=\max_{j\in N}P[\sigma^{1}]/P[\sigma^{j}]\leq 1/R$, and set $j^{*}=1$ if no such $j$ exists.
4  Calculate $k=\min_{i\in[j^{*}]}s_{k}=1$, and $k=1$ if no such index exists.
5else
6  Calculate $j^{*}=\max_{j\in N}P[\sigma^{1}]/P[\sigma^{j}]\leq R$, and set $j^{*}=1$ if no such $j$ exists.
7  Calculate $k=\min_{i\in[j^{*}]}s_{k}=0$, and $k=1$ if no such index exists.
8 end if
@@
Figure 1: Illustration of strategy in Algorithm 1 for firm, given $N=5$ candidates with $v_{1}>v_{2}=0$ and permutations given by Plackett-Luce. The x axis varies the $\gamma$ busy penalty, while the y axis varies accuracy as parameterized by the $1/\beta$ Gumbel noise parameter (higher values increase accuracy). Shades of red indicate regions where first-free selection is the optimal strategy and shades of blue indicate where first-busy selection is (darker shades indicate regions where the firm has larger $j^{*}$ or $k^{*}$ and thus “hunts” further down the list to find a candidate with their preferred free/busy status. The purple region is where the optimal strategy is “follow the ranking”. Note that the choice of first-free or first-busy depends only on $\gamma$, while the window size depends on the accuracy of the ranking tool.
@@
5 Superstar: societal welfare

### v3

@@ Abstract
Motivated by online platforms such as job markets, we study an agent choosing from a list of candidates, each with a hidden quality that determines match value. The agent observes only a noisy ranking of the candidates plus a binary signal that indicates whether each candidate is “free” or “busy”. Being busy is positively correlated with higher quality, but can also reduce value due to decreased availability. We study the agent’s optimal selection problem in the presence of ranking noise and free-busy signals and ask how the accuracy of the ranking tool impacts outcomes. In a setting with one high-valued candidate and an arbitrary number of low-valued candidates, we show that increased accuracy of the ranking tool can result in suboptimal social outcomes. For example, increased accuracy may mean that agents may be more likely to make offers to busy candidates, and (counter-intuitively) may be more likely to select lower-ranked candidates. We further discuss conditions under which these results extend to more general settings.
@@ 1 Introduction
Consider the following setting: you are a hiring manager, trying to find a candidate to interview for a hiring slot. You know that some candidates are better suited for your role than others - ideally, you would like to interview the best candidate available. To assist you with this goal, you might obtain a (noisy) ranking of the candidates from a commercial employment site, such as LinkedIn, Indeed, or other ranking tools. If the ranking is “good” (if the expected utility of candidates decreases as you navigate further down the list), the best action for the employer is clear: always pick the top-ranked candidate. Also, as the ranking tool becomes more accurate, the employer is only happier with their choice, as the expected value of the top-ranked candidate only increases.
However, in many cases, employers have access to side information, besides the ranking. For example, a hiring manager may know that some candidates are less likely to accept any eventual offer because they are already employed (see Filippas et al. (2023) for an empirical example of such a setting). We call such a candidate “busy”, as compared to a “free” candidate without these competing sources of employment. If a busy candidate would never accept an offer, then the choice is clear: the best option is to select the highest-ranked free candidate. However, if a busy candidate still has some chance of accepting an offer, the possible strategy space changes. In many markets there is an opposing force to the penalty for being busy: it may be the case that high-value candidates are more likely to be busy, and thus being busy is a signal of candidate quality. For example, this could be because high value candidates are more likely to have already been identified and hired by competing firms [^1] Throughout this paper, we will make the assumption that some candidates are better fits for the job than others (e.g. higher or lower value). We recognize that this assumption drastically simplifies the multi-faceted strengths and weaknesses each candidate may bring, and discuss this (and other) assumptions in more depth in Section 2.. In many settings it may not be possible to resolve the uncertainty inherent in a busy signal: for example, the candidate themselves may not know whether or not they would be willing to move without going through the interview process, at which point the firm has already used up its interview slot.
@@
Moreover, a firm’s hiring strategy has impacts on the rest of society. For example, consider that a firm’s choice to interview a busy (already employed) candidate over a free (unemployed) candidate may be more or less valuable to the candidate. An unemployed candidate may have higher utility for an interview than an employed candidate. Alternatively, an employed candidate may have a stronger ability to negotiate a better offer, which could increase their welfare more. While we do not directly model such externalities, we will use “probability of picking a busy candidate” and other such quantities as benchmarks to understand the impact of a ranking tool on broader society. Specifically, we will be interested in how these quantities change as the accuracy of the ranking tool increases. Do these measures uniformly improve or degrade as the firm uses a more accurate ranking tool?
@@
Motivation: More broadly, in this paper our goal is to study a stylized model of an agent using an algorithmic tool to make a decision, and combining the information encoded by the tool with her own source of information. Given the importance attached to continually increasing the accuracy of algorithmic tools, we wish to show that when these tools are used as part of a broader system involving human discretion and alternative sources of information, the returns to increased algorithmic accuracy are less clear: in particular, relevant measures of welfare could deteriorate under increased accuracy. We hope that our results will lead to further research in other scenarios where algorithmic tools are used alongside other sources of information, and in particular into what properties of such algorithmic tools are helpful for societal welfare. While we will use the employer hiring job candidates as our motivating example, we note that this setting is quite broad and could model a range of other scenarios where humans make decisions. For example, consider a hungry customer navigating a ranking of restaurants, each “crowded” or “not crowded”: a crowded restaurant is more likely to be good, but also has a longer wait.
@@
Summary: The rest of this paper is as follows: in Section 2, we will present our model and assumptions. In Section 3, we discuss the connection of our paper to related work, including the extensive literature on herding and the Pandora’s Box problem. Next, Section 4 analyzes the optimal strategy that a firm navigating a ranking might take: should she pick the first busy candidate, the first free candidate, or some other candidate entirely? In Section 5 we analyze phenomena related to welfare, like how frequently firms preferentially select job candidates who are already employed (potentially modeling unemployment discrimination) or avoid doing so (potentially modeling collusion in job hiring). We also analyze how often firms select the top ranked candidate, a measure of how incentive compatible the ranking is with potential implications for the welfare of the company creating the ranking tool. Throughout, we consider how increasing the accuracy of the ranking tool can affect the strategy, welfare, and fairness results in previous sections. Surprisingly, we show that increased accuracy can cause negative outcomes. In Section 6 we discuss extensions of our model and in Section 7 we conclude by discussing implications of our results. All proofs are deferred to the appendix.
@@ 2 Model, notation, and assumptions
Model and notation: First, we present our model. There are $N$ candidates, of which the agent (alternatively, firm) wishes to select exactly one: for example, to interview a job candidate. The candidates have different true values to the firm, where $v_{i}$ denotes the value of candidate $i$, though the firm cannot know the true value of candidates at each ranking (this only holds in the limit of a perfectly noiseless algorithmic ranking tool). The goal of the firm is to maximize their expected utility. In order to help with this decision, the firm has access to a ranking tool which produces noisy permutations over the candidates: $\sigma\sim\mathcal{P}$, where $\sigma_{i}$ denotes the index of the candidate ranked $i$th in permutation $\sigma$ with value $v_{\sigma_{i}}$. We will require that the expected value of candidates decreases in the ranking ($\mathbb{E}_{\sigma\sim\mathcal{P}}[v_{\sigma_{i}}]$ is decreasing in $i$).
@@
Permutation distribution: When comparing two distributions $\mathcal{P}$ and $\mathcal{P}^{\prime}$, we will use notation $P[\sigma]$ and $P[\sigma]^{\prime}$ to denote the probability $\mathcal{P}$ and $\mathcal{P}^{\prime}$ place on $\sigma$, respectively. In general, we will assume that the probability of observing each permutation is drawn from a Plackett-Luce distribution Plackett (1975); Luce and others (1959)). This commonly-used model is where i.i.d. noise is added to the true value of each item to produce $\hat{}v_{i}=v_{i}+\epsilon$, for $\epsilon$ drawn from a Gumbel distribution: then, the items are sorted in order of the noised values $\{\hat{}v_{i}\}$. Definition 3 defines how we will consider changing accuracy throughout most of the paper:
Definition 1 (Increased accuracy).
A permutation distribution governed by Plackett-Luce has increased accuracy when it has a smaller degree of noise (smaller standard deviation) in the Gumbel distribution.
However, many of our results rely only weakly on the Plackett-Luce model, and in the relevant appendices we state the sufficient properties on the permutation distribution in order for our results to apply.
@@
Firm strategy: In addition to the ranking, the firm has access to a single bit of side information about each candidate: specifically, a status vector $s$, where $s_{i}=1$ if the candidate ranked in position $i$ is free, and $s_{i}=0$ if the candidate is busy. We assume that higher-valued candidates are less likely to be available, in particular, that $v_{i}\geq v_{j}$ implies $p_{i}\leq p_{j}$, where $p_{i}$ gives the probability that candidate $i$ is free. We also assume that each candidate’s probability of being free or busy is independent of every other candidate: specifically, each free/busy realization is drawn from a Bernoulli distribution $s_{i}\sim Bern(p_{i})$ before the firm observes the ranking. Throughout this paper we will take these probabilities as fixed and assume each candidate reports her free-busy status honestly: in the conclusion we will discuss extensions of this model.
To summarize, there are three steps that occur in this model:
  1. 1.
    A permutation over candidates is drawn, with some $\sigma$ drawn from a Plackett-Luce distribution based on values of the candidates.
  2. 2.
    Each candidate obtains a free/busy status, independently given her probability $p_{i}$ of being free.
  3. 3.
    A firm may pick a single candidate, observing only the free-busy status and not the candidate values.
@@ 3 Related work
Our model has connections to celebrated models of agents strategically picking items, such as the hiring problem and Pandora’s box problem Krengel and Sucheston (1977); Weitzman (1978). However, our setting has key differences from these. Foremost among them is the signaling mechanism: rather than each item having an independent, known distribution of value, we model the scenario where the free/busy signal is generated by items with different probabilities, given different values. This structure, in conjunction with the ranking, means that the free/busy status of item $i$ can affect the expected value of item $j$ – generally not captured in existing models. Furthermore, in our setting the agent must pick a single item and commit to it without being able to observe its true value. This captures the setting where it is difficult or impossible to evaluate the item without committing to it first – for example, in hiring, when a candidate’s true “value” may not become apparent until after weeks or months of work.
@@
Additionally, our work has connections to long literature on herding (also known as information cascades) Banerjee (1992); Bikhchandani et al. (1992); Welch (1992), a pattern where it may be optimal for a decision-maker to follow actions taken by previous people, even ignoring their own information. Such patterns can lead to “cascades” where multiple sequential decision-makers each follow actions taken by previous decision-makers, rather than following their own information. In our setting, when an agent chooses to pick a lower ranked busy candidate over a higher ranked free, this could be seen as herding, since the free/busy status of candidates could be seen as created by actions taken by previous decision-makers, whereas the ranking is assumed to be private (personal) information. To our knowledge, herding has not been extensively studied in ranking settings: our model interestingly allows for herding to be moderated by the strength of private information through the quality of the ranking and the index of the first busy candidate.
@@
Our paper also has connections to the literature on algorithmic monoculture Kleinberg and Raghavan (2021), which studies when utility may be decreased when two agents rely on the same algorithmic ranking, rather than their own (uncorrelated) rankings. Variants of this setting have been explored, such as how algorithmic monoculture can lead to individuals experiencing correlated outcomes (“outcome homogenization”) Bommasani et al. (2022), or the implications of monoculture and of noisy matching in two-sided matching markets. For example, Peng and Garg (2023, 2024); Baek et al. (2025) all model settings where firms rely on common algorithmic tools to make hiring decisions, and the potential pitfalls of such an approach.
@@
Common features include the model of multiple agents taking strategic actions in relation to imperfect rankings, as well as the notion of a penalty for picking an item that another agent has picked. Key differences in our model include the fact that this penalty is softened – in particular, an agent can derive non-zero utility for picking an item that another has already chosen. This dramatically complicates the strategy space. Multiple works have considered the question of fairness in ranking - Zehlike et al. (2021) offers a helpful survey. In particular, Singh et al. (2021) describes a notion of fairness relating to how often items are presented in the top $k$, given that they have some probability of truly being the best $k$ items (relating to our notion of an item being “picked”). Issues of fairness and strategy have also been studied in the related area of algorithmic monoculture Black et al. (2022); Jain et al. (2023); Cooper et al. (2023); Hsu and Calmon (2022), while Peng et al. (2023) studies the tradeoff between diversity and accuracy, showing that consumption constraints (e.g. only top item can be consumed) explains away a tension between these two goals.
@@
Finally, there are multiple papers (empirical and theoretical) exploring models of firms using prediction tools to compete with each other. For example, Filippas et al. (2023) explicitly studies an empirical version of this where job candidates could pay to send a signal of availability to potential employers. Intriguingly, in this setting it seemed like such a signal did not lead employers to preferentially pick busy candidates, with most employers choosing a “first-free” strategy, perhaps indicating a parameter regime where this strategy was typically dominant. Horton et al. (2023) explores this setting in a dynamical pricing model, giving conditions where pricing levels for such a signal induce a truthful response. Additionally, Jagadeesan et al. (2023) explores a setting where improved data representation (as modeled by reduced Bayes risk) can paradoxically increase the error that users experience when they choose between two competing firms. Castera et al. (2022) studies statistical discrimination in stable matchings where candidates have preferences over firms, who only have noisy access to signals about candidate quality, and Castera et al. (2024) studies a setting where different candidates have differing degrees of correlation in their predictions across models, which leads to tensions between societal welfare and individual candidate welfare.
@@ 4 Superstar: Optimal Selection Strategies
Define $R=\frac{v_{1}/\gamma_{1}-v_{2}}{v_{1}-v_{2}/\gamma_{2}}\cdot r$.
if $R\leq 1$ then
  Calculate $j^{*}=\max_{j\in N}$ such that $P[\sigma^{1}]/P[\sigma^{j}]\leq 1/R$, and set $j^{*}=1$ if no such $j$ exists.
  Calculate $k=\min_{i\in[j^{*}]}$ such that $s_{k}=1$, and $k=1$ if no such index exists.
else
  Calculate $j^{*}=\max_{j\in N}$ such that $P[\sigma^{1}]/P[\sigma^{j}]\leq R$, and set $j^{*}=1$ if no such $j$ exists.
  Calculate $k=\min_{i\in[j^{*}]}$ such that $s_{k}=0$, and $k=1$ if no such index exists.
end if
@@
Figure 1: Illustration of strategy in Algorithm 1 for firm, given $N=5$ candidates with $v_{1}>v_{2}=0$ and permutations given by Plackett-Luce. The x axis varies the $\gamma$ busy penalty, while the y axis varies accuracy as parameterized by the $1/\beta$ Gumbel noise parameter (higher values increase accuracy). Shades of red indicate regions where first-free selection is the optimal strategy and shades of blue indicate where first-busy selection is optimal. Darker shades indicate regions where the firm has larger $j^{*}$ or $k^{*}$ and thus “hunts” further down the list to find a candidate with their preferred free/busy status. The purple region is where the optimal strategy is “follow the ranking” (pick the top ranked candidate always). Note that the choice of first-free or first-busy depends only on $\gamma$, while the window size depends on the accuracy of the ranking tool.
@@
5 Superstar: welfare implications